What fraction, reduced to its lowest terms, is halfway between 4/5 and 0.9 ?
a) 1/2 b) 2/3 c) 17/20 d) 6/7 e) 4 1/2

Answer by Theo(13342)

4/5 is the same as 8/10
.9 is the same as 9/10
halfway between would be 8.5/10 = 17/20

that would be selection C.

8/10 is the same as 16/20
9/10 is the same as 18/20

18/20-16/20 = 2/20 / 2 = 1/20
16/20 + 1/20 = 17/20
18/20 - 1/20 = 17/20